### v3

# DoF-Gaussian: Controllable Depth-of-Field for 3D Gaussian Splatting

Abstract
Recent advances in 3D Gaussian Splatting (3D-GS) have shown remarkable success in representing 3D scenes and generating high-quality, novel views in real-time. However, 3D-GS and its variants assume that input images are captured based on pinhole imaging and are fully in focus. This assumption limits their applicability, as real-world images often feature shallow depth-of-field (DoF). In this paper, we introduce DoF-Gaussian, a controllable depth-of-field method for 3D-GS. We develop a lens-based imaging model based on geometric optics principles to control DoF effects. To ensure accurate scene geometry, we incorporate depth priors adjusted per scene, and we apply defocus-to-focus adaptation to minimize the gap in the circle of confusion. We also introduce a synthetic dataset to assess refocusing capabilities and the model’s ability to learn precise lens parameters. Our framework is customizable and supports various interactive applications. Extensive experiments confirm the effectiveness of our method. Our project is available at https://dof-gaussian.github.io/.
![[Uncaptioned image]](x1.png)
1 Introduction
Depth-of-field (DoF) refers to the distance between the closest and farthest objects in a photo that appears acceptably sharp. In practice, photographers can control the DoF by adjusting the camera’s aperture size or focus distance to capture images with either a wide DoF (all-in-focus) or shallow DoF (defocused). The shallow DoF effect is an important technique in photography, as it draws the viewers’ attention to the focal region by blurring the surrounding areas. Meanwhile, novel view synthesis aims to create realistic images from novel viewpoints based on a set of source images. However, novel view synthesis typically requires all-in-focus input images and lacks the capability to render varied DoF effects, limiting its applications. In this work, we aim to render novel views with controllable DoF, adding cinematic quality to the results.
Images captured from the real world sometimes have shallow DoF. Specifically, points of light that do not lie on the focal plane are projected onto the sensor plane as blurred circles, referred to as the circle of confusion (CoC), and this introduces bokeh blur to the captured image. Most novel view synthesis methods experience performance degradation when processing shallow DoF input images. To address this issue, DoF-NeRF [55] and LensNeRF [17] introduce lens-based camera models in volume rendering to explicitly enable controllable DoF effects. Meanwhile, methods like Deblur-NeRF [26], DP-NeRF [19], and PDRF [33] employ sparse kernels to model defocus blur. However, implicit rendering approaches face significant challenges in training and rendering efficiency. In the realm of 3D Gaussian Splatting (3D-GS) [16], methods such as BAGS [34] and Deblurring 3DGS [18] propose blur estimation networks to model the blur kernels or scaling factors for defocus deblurring, but they cannot accommodate controllable DoF.
In this paper, we present DoF-Gaussian, an efficient framework for controllable DoF in 3D-GS, addressing the limitations of existing methods in handling shallow DoF inputs. This problem is non-trivial because 3D-GS and its variants [58, 23, 11, 54, 57] are typically based on a pinhole camera model that assumes all-in-focus inputs, meaning both foreground and background appear clear. When input images contain bokeh blur, it becomes challenging to accurately construct scene geometry and render depth maps. However, rendering images with different DoFs relies heavily on precise depth estimation. Additionally, our 3D-GS approach assumes an idealized CoC, which inevitably differs from the CoC seen in real photographs, posing further challenges for accurate defocus deblurring. Furthermore, existing datasets in this field are primarily borrowed from deblurring applications, limiting our evaluation to the model’s defocus deblurring capability and overlooking other aspects of controllable DoF effects.
To address these challenges, we present a new controllable DoF 3D-GS method with the following contributions: First, we employ a lens-based model instead of a pinhole-based one. We make the lens optical parameters, such as aperture size and focus distance, learnable to enable control over the depth of field. Second, we introduce per-scene adjustments for depth priors to minimize errors in scene geometry and depth relationships, ensuring accurate depth-of-field control. Third, we propose a defocus-to-focus adaptation strategy that focuses on learning the focal region after the learnable lens parameters have converged, compensating for differences in CoC. Finally, we present a synthetic dataset designed to comprehensively evaluate the model, including its refocusing capabilities and its ability to learn accurate aperture size and focus distance.
Benefiting from the lens-based imaging model, our approach enables controlled depth of field, allowing it to handle not only shallow DoF image inputs but also unlock a range of interactive and engaging applications, as depicted in Figure 1. For instance, users can render novel view images with varied DoF, refocus on custom datasets, and even adjust the shape of CoC, e.g., from circular to pentagonal or hexagonal CoC . Moreover, users can dynamically change the DoF while moving the camera or zooming, rendering videos with cinematic effects. Extensive experiments demonstrate that our method outperforms state-of-the-art depth-of-field and defocus deblurring methods.
2 Related Work
Depth-of-Field Rendering. Rendering DoF effects from a single all-in-focus image has been widely explored in previous work. Bokeh blur occurs when light is projected onto the image plane as a circular region, rather than as a point. The size of CoC is affected by the diameter of the aperture (aperture size) and the distance from the camera to the focal plane (focus distance). Photos captured with a small aperture usually present a wide DoF, i.e., all objects appear sharp. In contrast, as the aperture diameter increases, objects near the focal plane remain sharp, while those farther away become blurred with a larger CoC. Physically based methods [1, 20] rely on 3D scene geometry information and are time-consuming. Some methods use neural networks [13, 8, 38, 14, 12] trained end-to-end to obtain shallow DoF images. The DoF effects in some studies [35, 6, 56, 45, 48, 43, 61, 25, 36, 37] are controllable but usually require an extra disparity map. Wang et al. [46] combines neural fields with an expressive camera model to achieve all-in-focus reconstruction from an image stack. In this work, we show the connection between DoF rendering and novel view synthesis using a lens-based camera model.
Image Deblurring. Blur can generally be categorized into two main types: camera motion blur and defocus blur. Previous studies [26, 33, 34, 7, 18, 44, 24] attempt to address this issue. Our task of rendering all-in-focus and sharp novel view images from shallow DoF inputs can be considered as a form of defocus deblurring. By modeling the underlying physical principles, we can achieve defocus deblurring by simulating the circle-of-confusion. Furthermore, our method allows for adjustments to lens parameters, enabling applications such as refocusing—capabilities that deblurring approaches like BAGS [34] cannot achieve. Existing datasets in the 3D DoF field are often borrowed from the deblurring domain [26] or designed solely to assess defocus deblurring ability [55]. To support a wider range of applications and validate the accuracy of the learned lens model, we introduce a synthetic dataset for more comprehensive evaluations.
Novel View Synthesis. Novel view synthesis allows the rendering of unseen camera perspectives from 2D images and their corresponding camera poses. Recent advancements in synthesis can largely be attributed to Neural Radiance Field [28] and 3D Gaussian Splatting [16]. However, real-world images are often not clean and well-calibrated. Previous work [49, 62, 4, 29, 27, 31, 44, 54, 42] attempts to cope with non-ideal conditions, such as low light, camera motion, bokeh effects, and various types of image degradation. To recover a sharp scene from images with bokeh blur, methods like Deblur-NeRF [26], DP-NeRF [19], and PDRF [33] model defocus blur using sparse kernels. NeRFocus [50] and LensNeRF [17] achieve defocus effects by incorporating a lens-based camera model in volume rendering. While DoF-NeRF [55] is similar to our approach, it is limited by the training and rendering efficiency and quality of NeRF. BAGS [34] and Deblurring 3DGS [18] use blur estimation networks to model blur kernels or scaling factors but cannot achieve controllable DoF effects due to the absence of a lens imaging model. Other studies, such as [47, 15], focus on HDR tasks and consider depth-of-field simultaneously. RGS [7] addresses defocus blur by introducing an offset to the 2D covariance matrices of Gaussians. [30] introduces a ray tracing rendering algorithm for particle-based representations to enable many advanced techniques, such as the shallow depth-of-field effect. A concurrent work [51] also explores adjustable depth of field. Our method differs in its controllable DoF mechanism, with the introduction of depth priors to ensure accurate scene depth. In addition, our new defocus-to-focus adaptation strategy eliminates the need for the focus localization network proposed in their work. Besides, we introduce a synthetic dataset for comprehensive evaluation.
3 Preliminary
3D Gaussian Splatting represents a 3D scene as a mixture of anisotropic 3D Gaussians, where each Gaussian is characterized by a 3D covariance matrix and mean :
The covariance matrix holds physical meaning only when it is positive semi-definite. Therefore, to enable effective optimization via gradient descent, is decomposed into a scaling matrix and a rotation matrix , as:
To splat Gaussians from 3D space to a 2D plane, the view transformation matrix and the Jacobian matrix , which represents the affine approximation of the projective transformation, are utilized to obtain the covariance matrix in 2D space, as:
Subsequently, a point-based alpha-blending rendering can be performed to determine the color of each pixel:
where represents the color of each point, defined by spherical harmonics (SH) coefficients. The density is computed as the product of 2D Gaussians and a learnable point-wise opacity.
4 Method
We propose DoF-Gaussian, a controllable DoF approach for 3D-GS. The overall pipeline is illustrated in Fig. 2. In Section 4.1, we first develop a lens-based imaging model based on geometric optics principles to control DoF effects. Next, we employ the per-scene adjustment of depth priors to guide the correct scene geometry, as described in Section 4.2. Finally, in Section 4.3, we apply defocus-to-focus adaptation to minimize the inconsistencies in the CoC and enhance the defocus deblurring performance.
4.1 Lens-based Depth-of-Field Model
The physical principles underlying imaging and DoF have been extensively studied in the field of geometric optics [10, 32]. In an idealized optical system, the light emitted from a spatial point is projected onto a corresponding point in the image plane following the principles of pinhole imaging, as illustrated in Fig. 3 (). However, real-world cameras operate based on the lens model, as shown in Fig. 3 (). We show the two cases of in-focus and out-of-focus imaging, respectively. A spatial point , located at a distance from the lens, is projected on the imaging plane as a circular region referred to as the circle of confusion (CoC). The diameter of this region, , can be determined by the aperture parameter and focus distance , according to the following equation:
The focus distance primarily controls the depth position of the focal plane of the image, i.e., the depth of the sharp area, while the aperture parameter determines the extent of the bokeh effect. For the point located at the focus distance , the emitted rays directly converge through the lens to the corresponding point in the image plane, resulting in the absence of a circle of confusion (CoC). This also implies that .
We set focus distance and aperture parameter as learnable for each input image, and these parameters are continuously updated with the optimization of 3D-GS. By modeling the lens camera, the output , which includes bokeh blur effects, can be derived from the 3D-GS rasterizer rendering result using our CUDA-based Algorithm 1.
Ideally, the confuse function is represented by the indicator function . To achieve a smooth and natural DoF effect, we substitute it with a differentiable function, as proposed in Busam et al. [3]:
where defines the smoothness of confuse transition and represents the distance between two pixels.
We supervise the training by computing the reconstruction loss between the output image and the shallow DoF input image , where the loss function is a combination of the and a D-SSIM term, with set to :
For inference, we set the aperture size to so that we can render fully clear images for novel view synthesis.
4.2 Per-Scene Adjustment of Depth Priors
Unlike previous work that assumes input images are all-in-focus and fully clear, it becomes challenging to accurately reconstruct scene geometry and render depth maps for bokeh blurring input images. However, rendering images with varying depths of field relies on precise depth information.
To address this issue, we propose using depth prior as useful guidance for the accurate reconstruction of scene geometry. However, due to the presence of bokeh blur in shallow DoF inputs, directly employing a monocular depth network to predict depth maps as pseudo ground truth to regularize the Gaussians rendered depth map, as done in [5, 21], does not yield satisfactory results. Inspired by [53], we propose adapting per-scene depth priors through fine-tuning a monocular depth network based on the scene’s sparse reconstruction. Specifically, we utilize COLMAP [40, 41] to obtain a sparse Structure-from-Motion (SfM) point cloud and create a set of 3D Gaussians. Consequently, we can derive per-view sparse depth maps by projecting the 3D point clouds after multi-view stereo processing. Although the resulting depth map is sparse, it is robust and can serve as a supervisory signal for training the per-scene depth priors. To address the scale ambiguity inherent in acquired depth maps, we apply a silog loss [9] to fine-tune the depth network [39, 22], expressed as:
where is the scale factor and is the number of pixels in the image used for calculation. We apply the rasterization technique from Rade-GS [59], which enables the robust rendering of depth maps for 3D-GS scenes. We predict the depth maps from multi-view input images using the fine-tuned depth network and use them as depth priors to regularize . The depth loss can be described as follows:
As illustrated in Fig. 4, it is evident that the depth priors significantly enhance the geometric structure, resulting in a more accurate depth map compared to both BAGS [34] and the variant of our method without depth priors. In Section 5.3, we further conduct an ablation study comparing several other depth supervision strategies to demonstrate the effectiveness of our approach.
Method | Deblur-NeRF [26] | DoF-NeRF [55] | DP-NeRF [19] | BAGS [34] | Deblurring 3DGS [18] | Ours | ||||||||||||
PSNR | SSIM | LPIPS | PSNR | SSIM | LPIPS | PSNR | SSIM | LPIPS | PSNR | SSIM | LPIPS | PSNR | SSIM | LPIPS | PSNR | SSIM | LPIPS | |
Cake | 26.27 | 0.780 | 0.128 | 24.53 | 0.731 | 0.206 | 26.16 | 0.778 | 0.127 | 27.21 | 0.818 | 0.108 | 26.88 | 0.803 | 0.116 | 26.83 | 0.808 | 0.101 |
Caps | 23.87 | 0.713 | 0.161 | 22.67 | 0.636 | 0.261 | 23.95 | 0.712 | 0.143 | 24.16 | 0.725 | 0.159 | 24.50 | 0.742 | 0.149 | 24.62 | 0.749 | 0.147 |
Cisco | 20.83 | 0.727 | 0.087 | 20.64 | 0.724 | 0.107 | 20.73 | 0.726 | 0.084 | 20.79 | 0.743 | 0.070 | 20.83 | 0.732 | 0.079 | 21.00 | 0.744 | 0.069 |
Coral | 19.85 | 0.600 | 0.121 | 19.83 | 0.570 | 0.240 | 20.11 | 0.611 | 0.118 | 20.53 | 0.628 | 0.111 | 19.78 | 0.608 | 0.131 | 20.37 | 0.630 | 0.109 |
Cupcake | 22.26 | 0.722 | 0.116 | 21.89 | 0.706 | 0.143 | 22.80 | 0.741 | 0.096 | 22.93 | 0.762 | 0.080 | 22.11 | 0.734 | 0.099 | 22.97 | 0.757 | 0.079 |
Cups | 26.21 | 0.799 | 0.127 | 25.26 | 0.765 | 0.202 | 26.75 | 0.814 | 0.104 | 26.27 | 0.823 | 0.104 | 26.28 | 0.824 | 0.103 | 26.01 | 0.817 | 0.100 |
Daisy | 23.52 | 0.687 | 0.121 | 23.22 | 0.658 | 0.194 | 23.79 | 0.697 | 0.108 | 23.74 | 0.746 | 0.062 | 23.54 | 0.731 | 0.095 | 23.93 | 0.735 | 0.071 |
Sausage | 18.01 | 0.500 | 0.180 | 17.86 | 0.488 | 0.280 | 18.35 | 0.544 | 0.147 | 18.76 | 0.574 | 0.110 | 18.99 | 0.570 | 0.141 | 19.11 | 0.576 | 0.119 |
Seal | 26.04 | 0.777 | 0.105 | 24.85 | 0.687 | 0.143 | 25.95 | 0.778 | 0.103 | 26.52 | 0.812 | 0.090 | 26.18 | 0.817 | 0.098 | 26.57 | 0.825 | 0.087 |
Tools | 27.81 | 0.895 | 0.061 | 26.21 | 0.854 | 0.128 | 28.07 | 0.898 | 0.054 | 28.60 | 0.913 | 0.046 | 27.96 | 0.910 | 0.058 | 28.29 | 0.913 | 0.051 |
4.3 Defocus-to-Focus Adaptation
The CoC effect achieved under ideal optical imaging conditions, as described in Algorithm 1, inevitably differs from the real CoC effect produced by an actual DSLR camera. This discrepancy hinders our ability to accurately model the out-of-focus blur in shallow DoF input images.
To achieve sharper scenes, we propose an adaptation process that allows transitions from defocus to focus. Specifically, modeling the defocus effects with bokeh blur facilitates learning accurate lens parameters. Consequently, we place greater emphasis on capturing the sharp areas in training images once the lens parameters and have been optimized and converged. This approach is effective because, when the focus distance is determined, the sharp regions in the scene become approximately identifiable. After optimization iterations, we reweight the and loss by employing a step-like function:
where and are hyperparameters, and .
In addition, the aperture size is generally large after the convergence, which can cause blurring of pixels that should be in focus. Hence, we also reweight the aperture size on a pixel-wise basis:
Full Objective. We derive the final loss terms by incorporating the reconstruction loss from Eqn. 7, the depth loss from Eqn. 9, and an additional normal consistency loss as proposed by Huang et al. [11]. The normal consistency loss ensures that the Gaussian splats align with the surface by measuring the consistency between the normal directions computed from the Gaussian and the depth map:
where represents the surface normal direction obtained by applying finite differences on the depth map, indexes the intersected splats along the ray and denotes the blending weight of the intersection point.
Our final training loss is:
where is Hadamard product.
5 Experiments
Implementation Details. In our experiments, we set the smoothness of confuse transition . In addition, we set , , and for the defocus-to-focus adaptation. Our method is built upon Mip-Splatting [58], and our optimization strategy and hyperparameter settings remain consistent with it. We train each scene for iterations and set loss weights and .
Baselines and Evaluation Metrics. We compare our method with state-of-the-art deblurring techniques [34, 26, 18] and a depth-of-field method [55]. To quantitatively evaluate the quality of novel view images, we adopt widely used metrics such as PSNR, SSIM [52], and LPIPS [60]. We will compare with Wang et al. [51] once they release the code or provide the full experimental results in their paper.
Datasets. Following [34, 18, 26], we evaluate our model on the Deblur-NeRF dataset [26]. Subsequently, we conduct further evaluations on the Real Forward-facing dataset [28] and T&T_DB dataset [16]. The effectiveness of our method is validated on these datasets for both shallow DoF inputs and normal inputs. Furthermore, we also introduce a dataset in Section 5.1 to evaluate our model more comprehensively.
Method | PSNR | SSIM | LPIPS | ||
DoF-NeRF [55] | 25.59 | 0.788 | 0.207 | 0.196 | 0.256 |
Ours | 28.70 | 0.864 | 0.095 | 0.126 | 0.079 |
5.1 A Synthetic Dataset
We introduce a synthetic dataset, as illustrated in Fig. 6, based on depth estimation [2] and a single-image DoF rendering method [35] for each image in the Real Forward Facing dataset [28] and T&T_DB dataset [16]. This allows us to evaluate the refocusing ability of our model and to determine whether it accurately learns the correct aperture size and focus distance. In particular, by feeding the DoF rendering method [35] the known aperture size and focus distance, we can convert the all-in-focus images into shallow DoF images. Unlike existing defocus blur datasets proposed by Ma et al. [26] and Wu et al. [55], the test set in our dataset comprises images with shallow DoF. With known lens parameters, our model fits this shallow DoF effect to quantitatively measure the refocusing ability. Meanwhile, we establish the known lens parameters beforehand as ground truth, allowing us to calculate the errors and between the learned parameters and the ground truth. More details can be found in the supplementary material.
5.2 Comparisons
Quantitative Comparisons. We quantitatively compare DoF-Gaussian against other state-of-the-art methods using a real-world defocus blur dataset [26]. As shown in Table 1, our method outperforms other state-of-the-art baselines. These results suggest that our method achieves superior defocus deblurring performance and produces higher-quality novel view synthesis images.
Qualitative Comparisons. Visual qualitative comparisons are presented in Fig. 5. Our method surpasses all other methods in terms of image fidelity, generating novel view images that are more faithful to the ground truth images and exhibit less blur. Specifically, the bottle caps and shelves are the sharpest in our rendered images.
All-in-focus Inputs.
Method | PSNR | SSIM | LPIPS |
Mip-Splatting [58] | 27.05 | 0.893 | 0.115 |
Ours | 27.81 | 0.902 | 0.117 |
To further validate the effectiveness of our method in the all-in-focus setting, we design an experiment using wide DoF images as inputs. We conduct a comparison between our method and Mip-Splatting [58] on two all-in-focus datasets, the Real Forward Facing dataset [28] and the T&T_DB dataset [16]. As shown in Table 3, our method achieves comparable or even better performance to Mip-Splatting on the average metrics across the datasets. This demonstrates that our lens-based model not only excels with shallow DoF inputs but also achieves good results with general input images.
Our Synthetic Dataset. As shown in Table 2, our method achieves superior refocusing ability through controlled DoF rendering, producing rendered images with the highest quality on our synthetic dataset. Furthermore, our method learns more precise aperture size and focus distance than DoF-NeRF [55]. Since Deblur-NeRF [26], BAGS [34], and Deblurring 3DGS [18] do not incorporate a lens model, they cannot generate novel view images with a specific DoF. As a result, these methods cannot be evaluated on this dataset due to their lack of refocusing capability.
5.3 Ablation Study
As shown in Table 4 and Fig. 8, we conduct ablation studies to evaluate the effectiveness of our designs. The comparison between and indicates that the lens-based imaging model has a significant impact. This lens model not only facilitates defocus deblurring but also provides the ability to control the depth of field. Next, the defocus-to-focus adaptation () and the per-scene adjustment of depth priors () further enhance the scene geometry and improve the deblurring of details. Ultimately, the combination of all components yields the highest performance gain, leading to a dB increase in PSNR over the baseline model.
In addition, we conduct an ablation study on different depth supervision methods to demonstrate the superiority of our per-scene adjustment of depth priors, as shown in Table 5. Direct supervision using depth maps predicted by the depth network without scene-specific fine-tuning hinders the reconstruction quality of 3D-GS. Likewise, supervision based on sparse depth maps projected from a 3D point cloud fails to yield optimal results. We show visualizations of depth maps rendered by different depth supervision strategies in the supplementary material.
Lens | Depth | Defocus-to-focus | PSNR | SSIM | LPIPS | |
✗ | ✗ | ✗ | 21.31 | 0.636 | 0.239 | |
✓ | ✗ | ✗ | 23.05 | 0.728 | 0.109 | |
✓ | ✗ | ✓ | 23.59 | 0.742 | 0.104 | |
✓ | ✓ | ✗ | 23.42 | 0.738 | 0.098 | |
✓ | ✓ | ✓ | 23.97 | 0.756 | 0.093 |
Method | PSNR | SSIM | LPIPS |
No fine-tuned depth | 23.53 | 0.739 | 0.114 |
Sparse depth | 23.44 | 0.737 | 0.113 |
Ours | 23.97 | 0.756 | 0.093 |
5.4 Interactive Applications
Thanks to the lens-based model, we can not only represent 3D scenes from shallow DoF inputs but also render interactive DoF effects (see Fig. 7). Various cinematography styles can be achieved through different camera configurations, and users can input custom datasets, whether all-in-focus or bokeh images. By controlling aperture size and focus distance, we can render novel view images with different depths of field. In addition, we can modify the shape of the CoC, such as changing it from a circle to a hexagon. Users can also adjust the depth of field while moving the camera or zooming to create cinematic effects. Unlike [55], our GS framework significantly enhances the training and rendering efficiency, allowing for faster and more seamless interactions. The details on processing time will be discussed in the supplementary material.
6 Conclusion
While 3D Gaussian Splatting methods have achieved impressive performance on a wide range of reconstruction tasks, the applications on controllable DoF effects remain challenging and understudied. In this paper, we propose DoF-Gaussian, a controllable DoF method for 3D Gaussian Splatting. Specifically, we develop a lens-based model rather than pinhole imaging to overcome the limitations imposed by shallow DoF inputs for 3D-GS. Furthermore, we propose the per-scene depth adjustment and a defocus-to-focus adaptation to guarantee the performance of defocus deblurring. We also introduce a synthetic dataset for a more comprehensive evaluation. Thanks to the imaging principles, our method supports various interactive applications.
Future work. Inspired by our experiments, we found that modeling real-world physical imaging principles can enable our method not only to handle shallow DoF inputs but also to perform effectively on general inputs. This insight motivates us to explore a combination of non-ideal conditions, such as shallow DoF inputs, sparse views, and varied lighting environments, which are more aligned with casual photography in our daily lives.
Acknowledgement. This study is supported under the RIE2020 Industry Alignment Fund – Industry Collaboration Projects (IAF-ICP) Funding Initiative, as well as cash and in-kind contribution from the industry partner(s).
Supplementary Material
To see the dynamic effect of our method and visual comparisons, please refer to our supplementary video. This document includes the following contents:
Details of our synthetic dataset.
Correctness of the proposed dataset.
Color space
Details on all-in-focus experiments.
Details of ablation studies.
Processing time.
Limitations.
Appendix A Details of our synthetic dataset
To quantitatively evaluate the refocusing ability and assess whether models learn accurate lens parameters, we introduce a synthetic dataset based on Real Forward Facing dataset [28] and Tanks and Temples dataset [16]. Specifically, we apply a state-of-the-art depth estimation method [2] to generate disparity maps from input images. Subsequently, we employ a single-image DoF rendering method [35], feeding both the input images and disparity maps into the network to produce images with bokeh blur, as shown in Fig. 9. We choose [35] to synthesize shallow DoF images primarily because it is predominantly based on traditional physical renderer despite the incorporation of neural networks. The rendered circle of confusion (CoC) in this approach will not be significantly differ from the CoC produced by our lens-based physical imaging model. In addition, we excluded Drjohnson and Playroom, two indoor 360° scenes, due to significant monocular depth estimation errors of multi-view input images in indoor environments. At the same time, the inability to generate poses_bounds.npy files for the Train and Truck scenes prevents the evaluation of DoF-NeRF on these two scenes. We maintain these two scenes for comparisons with future 3D-GS methods. To assess whether the model learns the exact aperture size and focus distance for each input image, we set these parameters artificially in advance. For the focus distance we set three cases, , and , corresponding to focus on the background, mid-ground, and foreground, respectively. Recognizing that the aperture size is closely related to the image resolution, we here normalize it to to facilitate the calculation of the error. We consider two cases for aperture size: and . When we have optimized the 3D-GS scene, we get the learned focus distance and aperture size for each training image. Now, we can we can calculate the lens parameter error as:
where and indicate the preset aperture size and learned aperture size, respectively, and N means the number of training images. The smaller this error is, the more accurate our learned aperture size is. Similarly we use the following formula to calculate the focus distance error:
where and are the preset focus distance and learned focus distance. We use these two metrics to assess whether the model has learned the correct lens parameters. As demonstrated in Tables 6 and 7, our method outperforms DoF-NeRF [55] in both refocusing ability and the accurate estimation of lens parameters. Furthermore, as illustrated in Fig. 10, our method generates novel views that are more faithful to the ground-truth images.
Method | DoF-NeRF [55] | Ours | ||||
PSNR | SSIM | LPIPS | PSNR | SSIM | LPIPS | |
Fern | 24.80 | 0.736 | 0.200 | 28.41 | 0.867 | 0.098 |
Flower | 27.11 | 0.82 | 0.213 | 29.43 | 0.889 | 0.070 |
Fortress | 29.78 | 0.846 | 0.186 | 32.22 | 0.926 | 0.059 |
Horns | 24.23 | 0.812 | 0.235 | 27.64 | 0.863 | 0.122 |
Orchids | 19.99 | 0.608 | 0.213 | 21.54 | 0.659 | 0.165 |
Room | 26.55 | 0.842 | 0.198 | 32.16 | 0.933 | 0.071 |
Trex | 26.65 | 0.853 | 0.207 | 29.53 | 0.910 | 0.082 |
Train | — | — | — | 22.71 | 0.676 | 0.216 |
Truck | — | — | — | 21.09 | 0.675 | 0.312 |
Compared to the previous datasets proposed by Ma et al. [26] and Wu et al. [55], which evaluate defocus deblurring ability, our dataset is specifically designed to assess refocusing capabilities. Both the training and test sets in our synthetic dataset consist of shallow DoF images. We hope that this dataset will facilitate future work in this field.
Method | DoF-NeRF [55] | Ours | ||
Fern | 0.204 | 0.263 | 0.089 | 0.102 |
Flower | 0.127 | 0.280 | 0.091 | 0.074 |
Fortress | 0.156 | 0.299 | 0.187 | 0.021 |
Horns | 0.234 | 0.205 | 0.197 | 0.075 |
Orchids | 0.189 | 0.219 | 0.097 | 0.087 |
Room | 0.276 | 0.278 | 0.066 | 0.116 |
Trex | 0.189 | 0.251 | 0.154 | 0.079 |
Train | — | — | 0.225 | 0.113 |
Truck | — | — | 0.258 | 0.148 |
Appendix B Correctness of the proposed dataset
We validate the accuracy of the synthesis strategy using the BLB dataset, which comprises 500 test samples, each containing paired all-in-focus and defocus images. All-in-focus images are processed through our synthesis pipeline, and the resulting synthesized defocus images are compared with the ground truth to calculate PSNR and SSIM metrics. The High PSNR and SSIM values indicate that the synthesized bokeh is close to the real, thereby confirming the effectiveness of our synthesis strategy.
PSNR | SSIM | |
Ours | 43.30 | 0.9932 |
Appendix C Color space.
We apply a gamma transform on the input image to convert it from sRGB color space to linear color space. Subsequently, we simulate the circle-of-confusion within the linear color space. Finally, gamma correction is performed to convert the image from linear space back to sRGB space. The gamma value is . This process will be further emphasized in our revised version.
Appendix D Details on all-in-focus experiments
As shown in Table 9, we present the per-scene breakdown results of Real Forward-facing [28] and T&T_DB [16] datasets. These results align with the averaged results presented in the main text. Our method is built upon Mip-Splatting [58], a robust 3D-GS approach for all-in-focus inputs. Evidently, our method demonstrates superior performance compared to Mip-Splatting in most scenes. This indicates that our method can not only handle shallow DoF inputs, but also performs excellent under general input conditions, specially on Real Forward-facing dataset.
Method | Mip-Splatting [58] | Ours | ||||
PSNR | SSIM | LPIPS | PSNR | SSIM | LPIPS | |
Fern | 27.87 | 0.910 | 0.062 | 28.23 | 0.917 | 0.062 |
Flower | 25.55 | 0.868 | 0.101 | 27.81 | 0.910 | 0.059 |
Fortress | 27.85 | 0.913 | 0.072 | 32.57 | 0.951 | 0.032 |
Horns | 29.47 | 0.952 | 0.049 | 30.29 | 0.954 | 0.052 |
Leaves | 22.42 | 0.848 | 0.091 | 22.36 | 0.850 | 0.089 |
Orchids | 21.58 | 0.800 | 0.105 | 22.10 | 0.821 | 0.090 |
Room | 33.92 | 0.963 | 0.057 | 34.63 | 0.973 | 0.052 |
Trex | 28.67 | 0.951 | 0.056 | 30.29 | 0.961 | 0.042 |
Train | 21.89 | 0.821 | 0.190 | 21.15 | 0.791 | 0.251 |
Truck | 25.43 | 0.888 | 0.129 | 24.61 | 0.872 | 0.180 |
Playroom | 30.50 | 0.916 | 0.223 | 30.93 | 0.924 | 0.242 |
Drjohnson | 29.44 | 0.890 | 0.249 | 29.37 | 0.895 | 0.258 |
Appendix E Details of ablation studies
Method | baseline | w/o lens | w/o depth | w/o adaptation | sparse depth | w/o fine-tuned depth | ||||||||||||
PSNR | SSIM | LPIPS | PSNR | SSIM | LPIPS | PSNR | SSIM | LPIPS | PSNR | SSIM | LPIPS | PSNR | SSIM | LPIPS | PSNR | SSIM | LPIPS | |
cake | 24.15 | 0.710 | 0.216 | 25.69 | 0.782 | 0.131 | 26.51 | 0.800 | 0.116 | 26.09 | 0.794 | 0.107 | 26.43 | 0.795 | 0.121 | 26.41 | 0.792 | 0.133 |
caps | 21.24 | 0.559 | 0.332 | 23.57 | 0.713 | 0.148 | 24.41 | 0.741 | 0.145 | 24.12 | 0.737 | 0.142 | 24.52 | 0.742 | 0.164 | 24.52 | 0.745 | 0.157 |
cisco | 20.77 | 0.732 | 0.114 | 20.85 | 0.743 | 0.069 | 20.95 | 0.742 | 0.071 | 20.88 | 0.739 | 0.067 | 20.72 | 0.734 | 0.079 | 20.76 | 0.736 | 0.082 |
coral | 19.66 | 0.568 | 0.288 | 19.51 | 0.599 | 0.147 | 19.86 | 0.608 | 0.122 | 19.71 | 0.602 | 0.133 | 19.87 | 0.605 | 0.132 | 19.89 | 0.603 | 0.132 |
cupcake | 21.72 | 0.686 | 0.198 | 22.09 | 0.742 | 0.089 | 22.82 | 0.757 | 0.079 | 22.63 | 0.752 | 0.080 | 22.74 | 0.752 | 0.0087 | 22.81 | 0.752 | 0.086 |
cups | 24.29 | 0.749 | 0.223 | 25.89 | 0.814 | 0.100 | 25.91 | 0.818 | 0.114 | 26.06 | 0.820 | 0.086 | 25.34 | 0.800 | 0.115 | 25.63 | 0.804 | 0.117 |
daisy | 18.00 | 0.493 | 0.299 | 23.35 | 0.734 | 0.062 | 23.33 | 0.721 | 0.086 | 23.54 | 0.724 | 0.069 | 22.88 | 0.706 | 0.114 | 22.80 | 0.700 | 0.119 |
sausage | 17.45 | 0.461 | 0.284 | 17.99 | 0.515 | 0.169 | 18.47 | 0.536 | 0.151 | 18.29 | 0.529 | 0.156 | 18.18 | 0.531 | 0.172 | 18.55 | 0.550 | 0.153 |
seal | 20.71 | 0.561 | 0.288 | 24.34 | 0.744 | 0.114 | 25.54 | 0.790 | 0.105 | 25.34 | 0.781 | 0.088 | 26.17 | 0.805 | 0.095 | 26.10 | 0.804 | 0.097 |
tools | 25.09 | 0.845 | 0.152 | 27.17 | 0.898 | 0.056 | 28.09 | 0.911 | 0.051 | 27.53 | 0.902 | 0.052 | 27.57 | 0.900 | 0.061 | 27.82 | 0.902 | 0.059 |
In this section, we present detailed results of the ablation experiments in our main paper. In Table 11, we show the per-scene breakdown results of the ablation studies—baseline, w/o lens-based imaging model, w/o per-scene depth priors, w/o defocus-to-focus adaptation, sparse depth supervision, and no fine-tuned depth supervision. This indicates that each component of our system plays an important role in improving the image deblurring quality. In addition, we demonstrate the effectiveness of our approach by showing a visual comparison of the depth maps rendered by 3D-GS under different depth strategies, as shown in Fig. 11.
Appendix F Processing time.
As shown in Table 10, we recorded the processing time for both our method and other approaches on a single NVIDIA RTX A6000 GPU. For both Deblur-NeRF [26] and DoF-NeRF [55], we follow the specified training iterations outlined in the original papers, and calculate the training time. Due to the underlying NeRF-based framework, their average training time on the defocus deblurring dataset [26] is approximately 20 hours and 11 hours, respectively. Furthermore, their inference time is observed to be notably slow, achieving frame rates below 1 FPS. For the 3D-GS methods—BAGS [34], Deblurring 3DGS [18], and our method, we uniformly train for 30k iterations and record the training time and FPS. Although our method incorporates a lens imaging model, the training time is only slightly affected and it remains faster than BAGS. Benefit from the 3DGS framework, all GS-based methods can achieve fast rendering, obtaining FPS of approximately 360. In addition to training 3D Gaussian Splatting model, it takes about 3 minutes to fine-tune the depth network for per-scene depth priors.
Appendix G Discussion on depth supervision.
We employ per-scene adjustments of depth priors to guide the reconstruction and ensure the accurate scene geometry and rendered depth maps. The effectiveness of this approach is demonstrated by ablation experiments. However, depth maps predicted by the fine-tuned depth network are not entirely accurate, and using these as pseudo-gt to supervise the depth maps rendered by 3D-GS introduces a degree of noise. This residual noise may impact the precision of the final depth maps, particularly in scenes with complex geometry. We therefore use a strategy of gradual decay of the depth loss weight . In particular, we gradually decay this weight to of the initial value.
Appendix H Limitations
Our method may encounter limitations when the blur is view-consistent, such as in cases where the camera maintains a fixed focal point, i.e., focusing on a single target). Specifically, when the multi-view inputs all focus on the foreground, our method may struggle to recover clear background information. Consequently, a sharp scene can only be reconstructed if the input images contain both focused foreground and focused background elements. Addressing defocus deblurring under view-consistent conditions may be feasible through the integration of image priors, which we consider as a direction for future work.